OpenStudy (artemisxrpg):

The diameter of each wheel of a bicycle is 26 inches. If you are traveling at a speed of 35 miles per hour on this bicycle, through how many revolutions per min are the wheels turning?

OpenStudy (jack1):



OpenStudy (jack1):

first step @artemisxrpg , tell me the circumference please?

OpenStudy (artemisxrpg):

169pi?

OpenStudy (artemisxrpg):

wait

OpenStudy (artemisxrpg):

338pi

OpenStudy (jack1):

...no, Circumference = pi x D or 2 x pi x r, think you're working out area there

OpenStudy (artemisxrpg):

o, thought it was 2pi*r^2. So 26pi?

OpenStudy (jack1):

yep, which approx = 40.841 ish... now 2nd part, how many inches/ minute is 35 miles/ hour?

OpenStudy (artemisxrpg):

35 mi/hr * 1hr/60 min * 63360in/1mi = 38016 in/min

OpenStudy (jack1):

...check your calcs there, right units but wrong result, i got 36960 in/min...?

OpenStudy (jack1):

(63360 * 35) = 2,217,600 2,217,600/60 = 36,960, yeah?

OpenStudy (artemisxrpg):

Yea, I just messed up the calculation

OpenStudy (jack1):

all good, wasn't sure if it was me or u actually as i work with metrics, so inches and feet and miles are kinda foreign

OpenStudy (jack1):

back to the q now tho, so you're wheel is covering 36960 inches every minute and every time it covers 40.84 inches its completed 1 full revolution

OpenStudy (jack1):

so final equation is 36960 inches / minute x 1rev / 40.84071 inches = answer in revs / min

OpenStudy (artemisxrpg):

I got 904.98 rev/min. The answer in the book says it's 452.5 though

OpenStudy (jack1):

only thing i can think of is are you sure that it said diameter, not radius?

OpenStudy (jack1):

i got 904.98 also, the only way that's wrong and the answer is 452 (which is half of 904) is if the tyre size is doubled

OpenStudy (artemisxrpg):

Just checked the book, it's diameter

OpenStudy (jack1):

then the book's wrong dude, our calcs are solid hey.

OpenStudy (jack1):

will just get a ref to make sure, standby. @jim_thompson5910 or @bibby : do you have a sec to check our calcs?

OpenStudy (jack1):

dammit!, circumference of the tyre is 81.68 not 40, mah bad

OpenStudy (jack1):

26 x pi = 81.68, sorry

OpenStudy (artemisxrpg):

ok, I got the answer

OpenStudy (jack1):

36960 /81.68140899 =452.49, sorry man

OpenStudy (artemisxrpg):

It's cool. So for problems like this, I have to find the circumferences, convert it to the right unit if need it and divide that by the circumferences?

OpenStudy (jack1):

yep, sorry for the run around there man

OpenStudy (jack1):

yep, always get everything in the same units of measure first

OpenStudy (artemisxrpg):

alright, thanks for the help.